OpenStudy (anonymous):

factor z^3 - 3z^2 + 3z - 1

OpenStudy (anonymous):

either z(z2-3z+3)-1 or z3-3z(z-1)-1

hero (hero):

z^2(z - 3)+3z-1..... something is weird about this

OpenStudy (anonymous):

i got the same answer as you hero...... i know it is

hero (hero):

Are you sure you wrote it down correctly?

OpenStudy (anonymous):

Its a work packet :/

hero (hero):

Wait...let me see something real quick...hold on...

OpenStudy (anonymous):

k

OpenStudy (anonymous):

The next problem is the same way, x^3 + 15x^2 +75x + 125

OpenStudy (anonymous):

yes this is about factor and remainder theorem am gonna show my competence watch out for me now

hero (hero):

I think you have to re-write it like this: \[-3z^2 + 3z + z^3 -1\]

hero (hero):

Then factor as follows:

OpenStudy (anonymous):

hero you're amazing

OpenStudy (anonymous):

makes sense now <3

hero (hero):

the last two terms are difference of cubes

OpenStudy (anonymous):

you're a boss thanks

hero (hero):

Ooooh wait....lol

OpenStudy (anonymous):

yeah i know i caught your mistake im good now haha

hero (hero):

\[-3z(z-1)+(z-1)(z^2+z+1) = (z-1)(z^2-2z+1)\]

hero (hero):

There ya go

hero (hero):

Is that what you get too?

OpenStudy (anonymous):

i got the first half of what you last wrote as my answer

hero (hero):

Yes, but it isn't done unless you use distributive property. z - 1 is common to both

hero (hero):

The final answer I got is the correct answer.

hero (hero):

It may even be able to be simplified further...

hero (hero):

yep, it can...\[(z-1)(z-1)(z-1)\] Final Answer!

hero (hero):

BIZARRE!

OpenStudy (anonymous):

the question is really correct

OpenStudy (anonymous):

i think you made a mistake hero i dont see how you could distribute beyond my answer

hero (hero):

Yes, the question is really correct, and the complete factorization is a product of three binnomals

hero (hero):

I didn't make a mistake. You just don't understand that you need to use distributive property when factoring if you have two common binomials. Your answer is not a complete factorization

hero (hero):

I can prove that my factorization is correct

hero (hero):

And actually, it can even be simplified further to: \[(z-1)^3\]

hero (hero):

scroll down to the "expanded form": http://www.wolframalpha.com/input/?i=%28z-1%29%28z-1%29%28z-1%29

OpenStudy (anonymous):

hero is right but i got a new method see it now..

hero (hero):

A-grade....Post It!

OpenStudy (anonymous):

the question needs not t be written again.. ;et z^3-3z^2+3z-1

hero (hero):

Are you sure 'bout that?

hero (hero):

And what happened to the OP?

OpenStudy (anonymous):

−3z(z−1)+(z−1)(z^2+z+1)=(z−1)(z^2−2z+1) i still dont get how you did that

hero (hero):

\[ab + ac = a(b+c) \].....think about it: \[a = z-1; b = -3z; c = z^2+z+1 \] Ponder it for a while............. and don't forget to check wolfram alpha...

OpenStudy (anonymous):

first of all let f(z)=z^3-3z^2+3z-1 let us now find a number to replace z which would make the function 0, zero that number is 1 so, f(1)=1^3-3*1^2+3*1-1 this gives us 0, zero now, z=1 z-1=0 hence (z-1) is a factor now divide the original question by (z-1) and you would get z^2-2z+1 factorise this too by quadratic theorem and you would get (z-1)(z-1) and so the final answer is (z-1)(z-1)(z-1) please this is under factor and remainder theorem.....

OpenStudy (anonymous):

thats why it didnt make any sense you said you distributed but you just factored out the (z-1)

hero (hero):

I said I used "distributive property"

hero (hero):

But yeah, I factored out the z -1

hero (hero):

Well, at least you understand now...

OpenStudy (anonymous):

yup

hero (hero):

Grade A....If your method works so great, how come I still did a complete factorization and posted the answer before you did?

OpenStudy (anonymous):

It's not a comparing wingspan sizes contest, different people arrive at the answer in different ways.

hero (hero):

My argument is, he came up with his "answer" after I'd already posted the answer. In other words, he already knew what the factors were AFTER THE FACT....I gave him a medal, but still...

hero (hero):

Anyways, we can work on that next one...

hero (hero):

Unless you already have it.

OpenStudy (anonymous):

yeah it took me like 15 seconds

hero (hero):

Oh nice...What did you get?

OpenStudy (anonymous):

(x+5)^3

hero (hero):

Perfect!

hero (hero):

You're getting the hang of this. Your teacher will be proud

OpenStudy (anonymous):

hero dnt be annoyed at all' mathematics is never one way mathematicians use several means to arrive at the same answers

hero (hero):

Either way, you still posted your solution after the fact. Any reason why it took you so long?

hero (hero):

Almost 20 minutes after I posted my solution

hero (hero):

I'm not even including when I first started working on this. That's 20 minutes after I posted my solution only. So, if it really takes that long to do it your way, I'm good..